OpenStudy (dimples97):

Tom has a collection of 30 CDs and Nita has a collection of 18 CDs. Tom is adding 1 CD a month to his collection while Nita is adding 5 CDs a month to her collection. Find the number of months after which they will have the same number of CDs. Show your work. Someone please help me understand this and solve this?

OpenStudy (kropot72):

Let T be the number of cd's that Tom has. Let N be the number of cd's that Nita has. T = 30 + m N = 18 + 5m where m is the number of months. When T = N we can write 30 + m = 18 + 5m which when solved gives m = 3 months.

OpenStudy (dimples97):

Thankyou!!!

OpenStudy (kropot72):

You're welcome :)